# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ 3.2 Existence of FSE
We now examine conditions under which full sharing can be achieved in a PNE. For a profile to be an FSE, it must induce a quality profile that constitutes an ESE under enforced sharing. The uniqueness of the ESE implies that, if an FSE exists, both the creators’ content qualities and their sharing decisions are uniquely determined. Therefore, the remaining question is whether the provided revenue allocation is enough to induce such a profile in equilibrium.
@@
The result builds on a key structural property: when $\rho=1$, full sharing becomes a dominant strategy for every creator. Proposition 2 offers an intuition: As long as $\mathbf{x}$ is strictly positive, $\rho=1$ will surpass every possible threshold $\tau_{i}$, no matter the choice of $x_{i}$. This guarantees that any PNE must involve full sharing, and existence and uniqueness of the resulting quality profile then follow from Proposition 3. Extensions of Proposition 4 to other forms of $Q_{\text{AI}}$ are discussed in Appendix F.
Generalizing Proposition 4 beyond $\rho=1$ requires careful attention. Although Propositions 2 and 3 provide a tight characterization of the (decoupled) best responses in $s$ and $x$, the main challenge is ensuring that no combined deviation in both is beneficial. The following Theorem 5 offers a sufficient condition for a value of $\rho$ to induce an FSE.
@@
Fix any $\rho$, and let $\mathbf{x}^{\star}$ be the unique ESE of the subgame $G(\rho)$. For every $i\in[n]$, define $x_{i}^{\max}$ to be the unique positive value satisfying the equation $c_{i}(z)=\mu z^{\gamma}$. If $\rho>\max_{i\in[n]}\left\{\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathbf{x}^{\star}_{-i}\right\|}\right\}$, then $(\mathbf{x}^{\star},\mathbf{1})$ is the unique FSE of the subgame $G(\rho)$.
Proof Sketch of Theorem 5.
Fix some creator $i$. The proof relies on $x_{i}^{\max}$, a quality such that any choice of $x_{i}>x_{i}^{\max}$ will result in a strictly negative utility for $i$, regardless of its sharing level. When $\mathbf{s}_{-i}=\mathbf{1}_{-i}$, the threshold from Proposition 2 becomes $\frac{x_{i}}{x_{i}+(1+\alpha)\left\|\mathbf{x}_{-i}\right\|}$. This threshold is strictly increasing in $x_{i}$; thus, if $\rho>\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathbf{x}^{\star}_{-i}\right\|}$, any profitable deviation of $i$ must involve $s_{i}=1$. If this holds for every $i$, Proposition 3 guarantees no profitable deviations in $\mathbf{x}$ exist as well. ∎
Although the condition in Theorem 5 is theoretically meaningful, it is not directly constructive for implementation. The key challenge is that the quantity $\left\|\mathbf{x}^{\star}_{-i}\right\|$, which determines the right-hand side of the inequality, depends implicitly on the ESE $\mathbf{x}^{\star}$, which in turn is a function of $\rho$. Generally, it is not possible to obtain a closed-form expression for $\mathbf{x}^{\star}$ in terms of $\rho$, making the inequality difficult to evaluate. We defer to Appendix E an analysis under additional assumptions on the cost functions, which yields sharper insights into the existence and characterization of FSEs.
@@ 4 The Platform’s Revenue Maximization Problem
We now turn to the platform’s strategic problem of selecting the revenue-allocation parameter $\rho\in[0,1]$ that maximizes its retained revenue, subject to the constraint that creators respond by fully sharing their data. This induces a bi-level optimization problem: the platform acts as a Stackelberg leader, anticipating the equilibrium behavior of creators in response to its choice of $\rho$.
@@ 4.1 Formulation and Relaxation
Let $\mathcal{X}:[0,1]\to[0,\infty)^{n}$ denote the mapping from $\rho$ to its corresponding unique ESE, $\mathcal{X}(\rho)$. The platform’s optimization problem, which we term Platform Optimization, is given by:
@@ 4.2 Our Approach
2:initialize $\hat{\rho}\leftarrow 0$ and $\hat{U}\leftarrow 0$
3:for $t=1\dots\left\lfloor 1/\delta\right\rfloor$ do
4:     let $\rho\leftarrow\frac{t}{\left\lfloor 1/\delta\right\rfloor}$
5:     Find $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathbf{x}(\rho)\right\|_{2}\leq\delta$
6:     if $(\mathbf{x},\mathbf{1})$ is an $\eta$-FSE in $G(\rho)$ then
7:         if $U_{P}(\mathbf{x},\mathbf{1};f_{\rho})>\hat{U}$ then
8:              $\hat{\rho}\leftarrow\rho$ and $\hat{U}\leftarrow U_{P}(\mathbf{x},\mathbf{1};f_{\rho})$
@@
We propose a simple search-based procedure for computing an approximate solution of PO, illustrated in Algorithm 1. It receives as input an instance $\mathcal{I}$ and two parameters, $\delta$ and $\eta$. We carefully determine the values of $\delta$ and $\eta$ to reach a certain approximation level; we elaborate later in Theorem 6. Algorithm 1 performs a grid search over the interval $[0,1]$, evaluating a discrete set of allocations parameters at resolution $\delta$. Starting with initialized variables (Line 2), the algorithm iterates through candidate values $\rho=\nicefrac{{t}}{{\left\lfloor 1/\delta\right\rfloor}}$ (Lines 38). For each candidate, it approximates the corresponding ESE (Line 5) and verifies whether it constitutes an $\eta$-FSE (Line 6). If this stability criterion is met, it updates the current best solution whenever the resulting platform revenue exceeds the previous maximum (Line 8). The algorithm ultimately returns the best-performing $\rho$ across all valid candidates (Line 9). The next result establishes that the algorithm returns a near-optimal solution with provable guarantees on both approximation quality and runtime.
Theorem 6.
Fix any instance $\mathcal{I}$ with strongly convex cost functions. There exist constants $A,B$ such that for any $\varepsilon>0$, the execution of Algorithm 1 with $\delta=\frac{\varepsilon}{A}$ and $\eta=\frac{\varepsilon}{4}$ can be implemented in time $O\left(\frac{A^{2}B}{\varepsilon^{2}}\right)$. Moreover, its output $\hat{\rho}$ induces an $\varepsilon$-FSE and satisfies $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq\mathrm{OPT}-\varepsilon$.
We defer the specification of $A=A(\mathcal{I})$ and $B=B(\mathcal{I})$ to Appendix C, but remark that under mild regularity assumptions on the cost functions, they scale polynomially in the number of creators $n$.
Proof Sketch of Theorem 6.
@@
Fix any instance $\mathcal{I}$ and $\varepsilon>0$, and let $\delta,\eta$ denote the parameters in Theorem 6 that are passed to the algorithm. Next, focus on any feasible solution $\rho^{\star}$ to PO, and let $\rho$ denote the closest value to $\rho^{\star}$ the algorithm considers in Line 4. Notably, $\left|\rho-\rho^{\star}\right|\leq\delta$. Figure 1 illustrates how $\rho^{\star}$ is mapped to the unique ESE of $G(\rho^{\star})$, $\mathcal{X}(\rho^{\star})$. Similarly, $\rho$ is mapped to $\mathcal{X}(\rho)$, the unique ESE of $G(\rho)$.
Step 1: Beneficial deviations. We define the function $\phi$ such that $\phi(\mathbf{\tilde{x}},\tilde{\rho})$ is the maximal gain from any creator’s deviation from $(\mathbf{\tilde{x}},\mathbf{1})$ in $G(\tilde{\rho})$. Because $\rho^{\star}$ is feasible, $(\mathcal{X}(\rho^{\star}),\mathbf{1})$ is the FSE of $G(\rho^{\star})$; therefore, $\phi(\mathcal{X}(\rho^{\star}),\rho^{\star})=0$. Using technical lemmas, we show that $\phi$ is Lipschitz continuous and bound its Lipschitz constant. Additionally, we use the Implicit Function Theorem to bound the Lipschitz constant of $\mathcal{X}(\cdot)$, circumventing its lack of a closed-form expression. The blue region in Figure 1 (filled with horizontal lines) depicts the set of pairs $(\tilde{\mathbf{x}},\tilde{\rho})$ such that $\phi(\tilde{\mathbf{x}},\tilde{\rho})\leq\eta<\varepsilon$. A careful selection of $\delta$ implies that $\rho$ and $\rho^{\star}$ are close, and so are $\mathcal{X}(\rho)$ and $\mathcal{X}(\rho^{\star})$; therefore, $(\mathcal{X}(\rho),\rho)$ lies within the blue region, implying that the condition in Line 6 holds and $\rho$ is an $\varepsilon$-FSE of $G(\rho)$.
Step 2: Revenue approximation. Similarly, we show that the function $U_{P}(\tilde{\mathbf{x}},\mathbf{1};f_{\tilde{\rho}})$ is Lipschitz continuous and upper bound its Lipschitz constant. This ensures that there exists a neighborhood around $(\mathcal{X}(\rho^{\star}),\rho^{\star})$ with pairs $(\tilde{\mathbf{x}},\tilde{\rho})$ that obtain $\varepsilon$-close revenue. We let the orange region in Figure 1, filled with vertical lines, depict all such pairs. Picking $\delta$ appropriately ensures that $(\mathcal{X}(\rho),\rho)$ belongs to the orange region. Thus, in case $\rho^{\star}$ achieves a revenue arbitrarily close to OPT, at least one nearby point $\rho$ on the discrete grid yields a revenue within $\varepsilon$ of optimal.
@@
Step 3: Resolving output mismatch. Although we have demonstrated in Steps 1 and 2 that $\mathcal{X}(\rho)$ satisfies both the $\varepsilon$-FSE and revenue guarantees when $\rho$ lies sufficiently close to $\rho^{\star}$, the algorithm does not evaluate $\mathcal{X}(\rho)$ directly. Instead, it computes an approximate profile $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$ (recall Line 5). This creates a mismatch: Even if $\mathcal{X}(\rho)$ is a valid and near-optimal solution, the algorithm might overlook it because the nearby profile $\mathbf{x}$ fails one of the checks in Lines 6 and 7. To resolve this, let the green region (with a checkerboard pattern) in Figure 1 denote all pairs $(\tilde{\mathbf{x}},\rho)$ such that $\left\|\tilde{\mathbf{x}}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$. Unlike the orange and blue regions, the green region is a Euclidean ball. Note that $(\mathbf{x},\rho)$ is in this region for any arbitrary $\mathbf{x}$ that may have been found in Line 5. By choosing $\delta$ small enough so that the green ball is fully contained within both the blue and orange regions, we ensure that all such $\mathbf{x}$ exhibit the same stability and revenue guarantees as $\mathcal{X}(\rho)$, and vice versa. Thus, the algorithm safely evaluates $\mathbf{x}$ in place of $\mathcal{X}(\rho)$, and its output $\rho$ satisfies the desired approximation guarantees. ∎

### v2

@@ 3.2 Existence of FSE
We now examine conditions under which full sharing can arise in a PNE. For a profile to be an FSE, it must induce a quality profile that constitutes an ESE under enforced sharing. The uniqueness of the ESE implies that, if an FSE exists, both the creators’ content qualities and their sharing decisions are uniquely determined. Therefore, the remaining question is whether the provided revenue allocation is enough to induce such a profile in equilibrium.
@@
The result builds on a key structural property: when $\rho=1$, full sharing becomes a dominant strategy for every creator. Lemma˜2 offers intuition: as long as $\mathbf{x}$ is strictly positive, $\rho=1$ will surpass every possible threshold $\tau_{i}$, no matter the choice of $x_{i}$. This guarantees that any PNE must involve full sharing, and existence and uniqueness of the resulting quality profile then follow from Lemma˜3. Extensions of Proposition˜4 to other forms of $Q_{\text{AI}}$ are discussed in Section˜6.
Generalizing Proposition˜4 beyond $\rho=1$ requires careful attention. Although Lemmas˜2 and 3 provide a tight characterization of the (decoupled) best responses in $s$ and $x$, the main challenge is ensuring that no combined deviation in both is beneficial. The following Theorem˜5 offers a sufficient condition for a value of $\rho$ to induce an FSE.
@@
Fix any $\rho\in[0,1]$. For every $i\in[n]$, define $x_{i}^{\max}$ to be the unique positive value satisfying the equation $c_{i}(z)=\mu z^{\gamma}$. If $\rho>\max_{i\in[n]}\left\{\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}\right\}$, then $(\mathcal{X}(\rho),\mathbf{1})$ is the unique FSE of the subgame $G(\rho)$.
Proof Sketch of Theorem˜5.
Fix some creator $i$. The proof relies on $x_{i}^{\max}$, a quality such that any choice of $x_{i}>x_{i}^{\max}$ will result in a strictly negative utility for $i$, regardless of its sharing level. When $\mathbf{s}_{-i}=\mathbf{1}_{-i}$, the threshold from Lemma˜2 becomes $\frac{x_{i}}{x_{i}+(1+\alpha)\left\|\mathbf{x}_{-i}\right\|}$. This threshold is strictly increasing in $x_{i}$; thus, if $\rho>\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}$, any profitable deviation of $i$ must involve $s_{i}=1$. If this holds for every $i$, Lemma˜3 guarantees no profitable deviations in $\mathbf{x}$ exist as well. ∎
3.3 Analysis under Power Cost Functions
Although the condition in Theorem˜5 is theoretically meaningful, it is not constructive. The key challenge is that the quantity $\left\|\mathcal{X}_{-i}(\rho)\right\|$, determining the right-hand side of the inequality, depends on the ESE $\mathcal{X}(\rho)$, which in turn is a function of $\rho$. Generally, it is not possible to obtain a closed-form expression for $\mathcal{X}(\rho)$ in terms of $\rho$, making the inequality difficult to evaluate. To enable more precise analysis, we focus on a tractable yet flexible class of power cost functions, a standard structure in economic and game-theoretic models of effort and production [53, 36, 50].
Definition 3.
An instance is called $(n,a_{\min},a_{\max},\theta)$-power cost if it consists of $n$ creators, where each creator $i\in[n]$ has a cost function $c_{i}(x_{i})=a_{i}x_{i}^{\theta}$ with common exponent $\theta\in(1,2]$ and cost coefficients $a_{i}\in[a_{\min},a_{\max}]$.
Power cost functions enable more precise analysis. The next proposition reveals that as creator ecosystems grow large, inducing full sharing becomes increasingly tractable, requiring only modest revenue-sharing commitments from the platform.
Proposition 6.
For any $a_{\min},a_{\max},\theta$ and $w\in(0,1]$, there exists $N(w)\in\mathbb{N}$ such that for all $n>N(w)$, in any $(n,a_{\min},a_{\max},\theta)$-power cost instance, the subgame $G(\rho)$ admits an FSE for every $\rho\in(w,1]$.
The proof of Proposition˜6 leverages previous work [50, Theorem 3] to construct a lower bound on $\left\|\mathcal{X}_{-i}(\rho)\right\|$ which depends on the number of creators $n$. We then obtain the desired result by substituting this lower bound in Theorem˜5 and extracting the necessary $n$.
While Proposition˜6 suggests that platforms can achieve FSE with low revenue sharing, this raises a natural question about content production incentives. Lower allocation parameters may reduce creators’ motivation to invest in quality, potentially decreasing overall platform traffic. The next proposition examines this relationship.
Proposition 7.
Fix any homogeneous power cost instance, i.e., $(n,a,a,\theta)$ for some $a>0$. Then, $\left\|\mathcal{X}(\rho)\right\|$ is strictly increasing in $\rho$.
Proposition˜6 and Proposition˜7 reveal competing incentives in the choice of $\rho$: retaining a larger share of existing GenAI revenue versus increasing total revenue through larger traffic. The next section navigates this tradeoff algorithmically.
@@ 4 The Platform’s Revenue Maximization Problem
We now turn to the platform’s strategic problem of selecting the revenue-allocation parameter $\rho\in[0,1]$ that maximizes its retained revenue, subject to the constraint that creators respond by fully sharing their data. This induces a bilevel optimization problem: the platform acts as a Stackelberg leader, anticipating the equilibrium behavior of creators in response to its choice of $\rho$.
@@ 4.1 Formulation and Relaxation
The platform’s optimization problem, which we term Platform Optimization, is given by:
@@ 4.2 Algorithm
[…]
4:for $t\leftarrow 1$ to $\lceil 1/\delta\rceil$ do
5:  Let $\rho\leftarrow\min\left\{\dfrac{t}{\lceil 1/\delta\rceil},1\right\}$
6:  Find $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$
7:  if $(\mathbf{x},\mathbf{1})$ is an $\eta$-FSE in $G(\rho)$ and $U_{P}(\mathbf{x},\mathbf{1};f_{\rho})>\hat{U}$ then
8:   $\hat{U}\leftarrow U_{P}(\mathbf{x},\mathbf{1};f_{\rho}),\hat{\rho}\leftarrow\rho$
@@
We propose a simple search-based procedure for computing an approximate solution of PO, illustrated in Algorithm˜1. It receives as input an instance $\mathcal{I}$ and two parameters, $\delta$ and $\eta$. We carefully determine the values of $\delta$ and $\eta$ to reach a certain approximation level; we elaborate later in Theorem 8. Algorithm˜1 performs a grid search over the interval $[0,1]$, evaluating a discrete set of allocations parameters at resolution $\delta$. Starting with initialized variables (Line 3), the algorithm iterates through candidate values $\rho=\nicefrac{{t}}{{\left\lfloor 1/\delta\right\rfloor}}$ (Lines 48). For each candidate, it approximates the corresponding ESE (Line 6) and verifies whether it constitutes an $\eta$-FSE (Line 7). If this stability criterion is met, it updates the current best solution whenever the resulting platform revenue exceeds the previous maximum (Line 8). The algorithm ultimately returns the best-performing $\rho$ across all valid candidates (Line 9). The next result establishes that the algorithm returns a near-optimal solution with provable guarantees on both approximation quality and runtime.
Theorem 8.
Fix any instance $\mathcal{I}$ with strongly convex cost functions. There exist constants $A,B$ such that for any $\varepsilon>0$, the execution of Algorithm˜1 with $\delta=\frac{\varepsilon}{A}$ and $\eta=\frac{\varepsilon}{4}$ can be implemented in time $O\left(\frac{A^{2}B}{\varepsilon^{2}}\right)$. Moreover, its output $\hat{\rho}$ induces an $\varepsilon$-FSE and satisfies $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq\mathrm{OPT}-\varepsilon$.
The exact specification of $A=A(\mathcal{I})$ and $B=B(\mathcal{I})$ involves somewhat technical details; however, both constants depend on upper and lower bounds of the equilibrium quality $\left\|\mathcal{X}(\rho)\right\|$, which under mild regularity assumptions are themselves polynomial in the number of creators $n$.
4.3 Proof Sketch of Theorem˜8
@@
Fix any instance $\mathcal{I}$ and $\varepsilon>0$, and let $\delta,\eta$ denote the parameters in Theorem˜8 that are passed to the algorithm. Next, focus on any feasible solution $\rho^{\star}$ to PO, and let $\rho$ denote the closest value to $\rho^{\star}$ the algorithm considers in Line 5. Notably, $\left|\rho-\rho^{\star}\right|\leq\delta$. Figure 1 illustrates how $\rho^{\star}$ is mapped to the unique ESE of $G(\rho^{\star})$, $\mathcal{X}(\rho^{\star})$. Similarly, $\rho$ is mapped to $\mathcal{X}(\rho)$, the unique ESE of $G(\rho)$.
Step 1: Beneficial deviations
[…]
Step 2: Revenue approximation
[…]
Step 3: Resolving output mismatch
@@
Although we have demonstrated in Steps 1 and 2 that $\mathcal{X}(\rho)$ satisfies both the $\varepsilon$-FSE and revenue guarantees when $\rho$ lies sufficiently close to $\rho^{\star}$, the algorithm does not evaluate $\mathcal{X}(\rho)$ directly. Instead, it computes an approximate profile $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$ (recall Line 6). This creates a mismatch: even if $\mathcal{X}(\rho)$ is a valid and near-optimal solution, the algorithm might overlook it because the nearby profile $\mathbf{x}$ fails one of the checks in Line 7. To resolve this, let the green region (with a checkerboard pattern) in Figure˜1 denote all pairs $(\tilde{\mathbf{x}},\rho)$ such that $\left\|\tilde{\mathbf{x}}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$. Unlike the orange and blue regions, the green region is a Euclidean ball. Note that $(\mathbf{x},\rho)$ is in this region for any arbitrary $\mathbf{x}$ that may have been found in Line 6. By choosing $\delta$ small enough so that the green ball is fully contained within both the blue and orange regions, we ensure that all such $\mathbf{x}$ exhibit the same stability and revenue guarantees as $\mathcal{X}(\rho)$, and vice versa. Thus, the algorithm safely evaluates $\mathbf{x}$ in place of $\mathcal{X}(\rho)$, and its output $\rho$ satisfies the desired approximation guarantees.
Efficient computation
First, we prove that the enforced sharing game belongs to the class of strongly monotone games [35]. This property allows us to apply multi-agent mirror descent to find the approximate quality profile required in Line 6 in $O(1/\delta)$ iterations [12], each requiring $O(n)$ time to evaluate creator gradients. The $\delta$-FSE condition can also be efficiently verified. For each creator, it suffices to solve a single one-dimensional optimization: maximize utility over $x_{i}\geq 0$ with $s_{i}=0$, holding the strategies of all other creators fixed. This follows from Lemma˜2, which ensures that any profitable deviation is dominated by a deviation that has $s_{i}=1$ or $s_{i}=0$. The former is implicitly covered by the approximation guarantee of the computed profile $\mathbf{x}$, and the latter can be efficiently approximated, as we demonstrate that the utility function is Lipschitz continuous under each fixed value of $s$. We choose $\delta^{\prime}$ small enough so that approximation errors from the ESE computation and the deviation check are jointly bounded by $\delta$. After finding the maximal utility under $s_{i}=0$, it remains only to test whether the resulting gain exceeds $\delta$, thereby completing the verification.
4.4 Refinements for Power Cost Functions
Having established the correctness and efficiency of Algorithm˜1, we now examine its behavior under specific cost structures. For the class of power cost functions (recall Definition˜3), we obtain explicit bounds on the algorithm’s runtime dependence on the number of creators.
Proposition 9.
Consider any $(n,a_{\min},a_{\max},\theta)$-power cost instance $\mathcal{I}$ with $a_{\min},a_{\max}=O(1)$. Then the constants $A(\mathcal{I})$ and $B(\mathcal{I})$ from Theorem˜8 satisfy $A(\mathcal{I})=O(n^{2})$ and $B(\mathcal{I})=O(n)$.
Although Algorithm˜1 offers an efficient method for achieving an anytime approximation of $\mathrm{OPT}$, its runtime inevitably depends on $n$. Thus, for large platforms, it may be preferable to use a fixed allocation rule rather than find one algorithmically, as the latter option might be impractical. It turns out that for many power cost instances, there is a closed-form choice of $\rho$ that induces an FSE and provides a multiplicative approximation of the platform’s optimal revenue. Interestingly, this choice of $\rho$ does not depend on $n$ or the coefficients $\left(a_{i}\right)_{i=1}^{n}$.
Proposition 10.
Consider large enough $n,\mu$ and any power cost instance such that $(\alpha+1)\gamma>\theta$. Then, choosing $\rho=\frac{\alpha\gamma+\gamma-\theta}{\alpha\theta}$ ensures that $\rho$ induces an FSE, and yields platform revenue satisfying $U_{P}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})\geq(2\alpha+2)^{-\frac{\gamma}{\theta-\gamma}}\cdot\mathrm{OPT}$.
Notably, larger values of $\theta$ decrease the expression of $\rho$ in Proposition˜10, and, perhaps counterintuitively, larger values of $\alpha$ increase it. We detail more about these phenomena in Section˜5, where they also arise from simulations on the optimal allocation parameter. The proof of Proposition˜10 builds on a lower bound of $\left\|\mathcal{X}(\rho)\right\|$ from [50], which we adapt to our setting. We then derive the closed-form expression for $\rho$ by substituting this lower bound into the platform’s revenue function and optimizing it analytically. Proposition˜6 ensures that the resulting profile is a valid FSE, while the approximation factor reflects the gap between the bound and the true value of $\left\|\mathcal{X}(\rho)\right\|$.